# Centralization vs. decentralization in multi-robot sweep coverage with ground robots and UAVs

### v2

@@ 5.2 Fault tolerance
We assess fault tolerance in terms of the impact that robot failures have on coverage completeness and uniformity, and sweep completion time. We run the fault tolerance experiments in the setup with 100 obstacles and, at time step 400, we impose failure on either two, four, or six of the eight total ground robots. When a ground robot fails, it stops moving but it continues to be included in the calculation of coverage uniformity. Thus, a failed ground robot stops improving coverage completeness and meanwhile consistently worsens coverage uniformity.
(a) Scalability: Coverage completeness.
(b) Scalability: Coverage uniformity.
(c) Scalability: Sweep completion time.
@@ 5.3 Scalability
As the number of robots increases, the coverage completeness achieved by the predetermined and formation approaches stays the same (perfect or nearly perfect). At the same time, the completeness achieved by the decentralized approaches improves rapidly, nearing the perfect completeness of the other approaches in the system size of 16 robots (97.7% to 98% completeness). The sweep completion time of all approaches also improves as system size grows, with the predetermined and formation approaches improving at a slightly faster rate. The predetermined and formation approaches also improve in terms of coverage uniformity as the system size increases, however, the coverage uniformity of the decentralized approaches, by contrast, steadily worsens. This suggests that the better completeness and faster speeds of the decentralized approaches are achieved through increased redundancy.
@@ 6 Discussion
Regarding fault tolerance, while the decentralized approaches achieved better coverage completeness under robot failures at the end of a sweep, this advantage disappears if completeness is considered in terms of elapsed time instead of sweep end. Also, the decentralized approaches never achieved better uniformity at the end of a sweep. Overall, the predetermined and formation approaches worsened at a faster rate, but their worsening mostly brought them down to a performance level close to that of the decentralized approaches. Given that the predetermined and formation approaches worsened at a faster rate, it is possible that they would under-perform the decentralized approaches in even worse conditions (i.e., worse than the 75% failure rate tested here). However, it is important to note that if the predetermined and formation approaches were permitted to adapt their sweep trajectories during operation according to the number of remaining robots after failure, their fault tolerance would substantially improve. In short, we conclude that fault tolerance through redundancy, although present, did not give the decentralized approaches a meaningful advantage over the the other approaches in the tested conditions. Of course, it is also important to note that the failure of UAVs was not tested, and that the predetermined, formation, and decentralized with beacons approaches would all be susceptible to this failure type, while the random walk approach would not.
@@ Acknowledgements
Authors’ contributions: AJ, MW, MA, MD, and MKH conceived and planned the experiments. AJ, MA, and WZ developed the code used in the experiments, supervised by MW and MA. AJ conducted the experiments and collected the data. AJ, MW, MD, and MKH led the interpretation of the experimental results and conducted the analysis. The manuscript was written by AJ, MW, MD, and MKH, led by MKH. All authors read and approved the final manuscript.
@@
The code used in the experiments is open-source, available on GitHub: https://github.com/BlueDiamond07/MultiRobotSweepCoverage.
@@
Algorithm 5 Predetermined: ground robot (more detailed version).
@@
Table 2: Sweep coverage performance in all three obstacle setups: coverage completeness, in % ($\alpha$); coverage uniformity ($\rho$); and sweep completion time, in time steps ($T$). The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 3: Sweep coverage performance in each obstacle setup, separately: coverage completeness, in % ($\alpha$); coverage uniformity ($\rho$); and sweep completion time, in time steps ($T$). The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 4: Fault tolerance results: coverage completeness, in % ($\alpha$); coverage uniformity ($\rho$); and sweep completion time, in time steps ($T$). The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 5: Fault tolerance results: change ($\Delta$) in mean (M) for coverage completeness, in % ($\alpha$); coverage uniformity ($\rho$); and sweep completion time, in time steps ($T$).
Change in mean for 0-2 failures
@@
Change in mean for 0-4 failures
@@
Change in mean for 0-6 failures
@@
Table 6: Scalability results: coverage completeness, in % ($\alpha$); coverage uniformity ($\rho$); and sweep completion time, in time steps ($T$). The table reports the mean (M) for all three metrics, as well as the standard deviation (SD), 95% confidence interval (CI), and mean (M) per robot for coverage completeness and uniformity.
@@
Table 7: Scalability results: change ($\Delta$) in mean (M) for coverage completeness and coverage completeness per robot, in % ($\alpha$); coverage uniformity and coverage uniformity per robot ($\rho$); and sweep completion time, in time steps ($T$).

### v4

@@ 5.2 Fault tolerance
We assess fault tolerance in terms of the impact that robot failures have on coverage completeness and uniformity, and sweep completion time. We run the fault tolerance experiments in the setup with 100 obstacles and, at time step 400, impose failure on either two, four, or six of the eight total ground robots. When a ground robot fails, it stops moving, but it continues to be included in the calculation of coverage uniformity. Thus, a failed ground robot stops improving coverage completeness and meanwhile consistently worsens coverage uniformity.
@@ 5.3 Scalability
As the number of robots increases, the coverage completeness achieved by the predetermined and formation approaches stays the same (perfect or nearly perfect). At the same time, the completeness achieved by the decentralized approaches improves rapidly, nearing the perfect completeness of the other approaches in the system size of 16 robots (97.7% to 98% completeness). The sweep completion time of all approaches also improves as system size grows, with the predetermined and formation approaches improving at a slightly faster rate. The predetermined and formation approaches also improve in terms of coverage uniformity as the system size increases. The coverage uniformity of the decentralized approaches, by contrast, steadily worsens. This suggests that the better completeness and faster speeds of the decentralized approaches are achieved through increased redundancy.
@@ 6 Discussion
Regarding fault tolerance, while the decentralized approaches achieved better coverage completeness under robot failures at the end of a sweep, this advantage disappears if completeness is considered in terms of elapsed time instead of sweep end. Also, the decentralized approaches never achieved better uniformity at the end of a sweep. Overall, the predetermined and formation approaches worsened at a faster rate, but their worsening mostly brought them down to a performance level close to that of the decentralized approaches. Given that the predetermined and formation approaches worsened at a faster rate, it is possible that they would under-perform the decentralized approaches in even worse conditions (i.e., worse than the 75% failure rate tested here). However, it is important to note that if the predetermined and formation approaches were permitted to adapt their sweep trajectories during operation according to the number of remaining robots after failure, their fault tolerance would substantially improve. In short, we conclude that fault tolerance through redundancy, although present, did not give the decentralized approaches a meaningful advantage over the other approaches in the tested conditions. Of course, it is also important to note that the failure of UAVs was not tested, and that the predetermined, formation, and decentralized with beacons approaches would all be susceptible to this failure type, while the random walk approach would not.
@@ 6.1 Future work
7 Conclusion
In terms of coverage completeness, coverage uniformity, and completion time, the more centralized approaches outperformed the more decentralized approaches and the predetermined approach outperformed all other approaches — as expected. The performance of our random billiards approach showed that it is adequately representative of the random walk category in the literature, for the sweep coverage task. Our decentralized approach with beacons performed roughly equivalently to our random billiards approach, suggesting that greater centralization would need to be integrated to start seeing performance benefits.
In terms of fault tolerance, scalability, and flexibility, although the decentralized approaches outperformed the more centralized approaches by some metrics, they did not display the strong outperformance that had been expected. For example, the decentralized approaches achieved better coverage completeness under robot failures at the end of a sweep, but not at the end of the same elapsed time. In terms of coverage uniformity, although the performance of the more centralized approaches worsened at a much faster rate as more robots failed, they still always outperformed the decentralized approaches in the tested conditions. As another example, in terms of scalability, the coverage completeness of the decentralized approaches improved drastically as the number of robots increased, but still never outperformed the more centralized approaches, because the more centralized approaches always had near-perfect coverage completeness regardless of the number of robots. Also, the coverage uniformity of the more centralized approaches improved as the number of robots increased, while that of the decentralized approaches worsened. Finally, regarding flexibility in response to more cluttered environments, there was no notable difference between the decentralized and the more centralized approaches, in the tested conditions.
Overall, the more centralized approaches greatly outperformed the decentralized ones in terms of sweep coverage performance (coverage completeness, coverage uniformity, completion time). The decentralized approaches showed much better fault tolerance as well as better scalability by some metrics (i.e., slower overall performance degradation under increasing robot failures, and faster coverage completeness improvement under an increasing number of robots), but these fault tolerance and scalability benefits did not cause the decentralized approaches to outperform the more centralized ones — rather, the performance of the decentralized approaches became more similar to that of the centralized ones, in the tested conditions. However, the conclusions of this study have been demonstrated only for a specific set of experimental conditions. There are many additional conditions that should be studied in future work, such as communication bottlenecks, failure of a supervisor robot, more complex environments, and scalability to much larger system sizes.
@@ Acknowledgements
Authors’ contributions: AJ, MW, MA, MD, and MKH conceived and planned the experiments. AJ, MA, and WZ developed the code used in the experiments, supervised by MW, MA, and MKH. AJ conducted the experiments and collected the data. AJ, MW, MD, and MKH led the interpretation of the experimental results and conducted the analysis. The manuscript was written by AJ, MW, MD, and MKH, led by MKH. All authors read and approved the final manuscript.
Conflicts of interest
The authors declare that they have no conflict of interest.
@@
The code used in the experiments is open-source, available on GitHub: https://github.com/BlueDiamond07/MultiRobotSweepCoverage. It is also available on Zenodo: https://doi.org/10.5281/zenodo.14846074.
@@
Algorithm 5 – Ground robot: Predetermined (more detailed version).
@@
Table 2: Sweep coverage performance in all three obstacle setups: coverage completeness ($\alpha$), in %; coverage uniformity ($\rho$); and sweep completion time ($T$), in time steps. The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 3: Sweep coverage performance in each separate obstacle setup (100, 200, and 300 obstacles): coverage completeness ($\alpha$), in %; coverage uniformity ($\rho$); and sweep completion time ($T$), in time steps. The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 4: Fault tolerance results: coverage completeness ($\alpha$), in %; coverage uniformity ($\rho$); and sweep completion time ($T$), in time steps. The table reports the mean (M) and standard deviation (SD) for all three metrics, as well as the 95% confidence interval (CI) for coverage completeness and uniformity.
@@
Table 5: Fault tolerance results for different failure rates compared to no failure: change ($\Delta$) in mean (M) w.r.t. 0 failures for coverage completeness ($\alpha$), in %; coverage uniformity ($\rho$); and sweep completion time ($T$), in time steps.
Change in mean for 2 failures
@@
Change in mean for 4 failures
@@
Change in mean for 6 failures
@@
Table 6: Scalability results: coverage completeness ($\alpha$), in %; coverage uniformity ($\rho$); and sweep completion time ($T$), in time steps. The table reports the mean (M) for all three metrics, as well as the standard deviation (SD), 95% confidence interval (CI), and mean per robot for coverage completeness and uniformity.
@@
Table 7: Scalability results: change ($\Delta$) in mean (M) for coverage completeness ($\alpha$) and coverage completeness per robot, in %; coverage uniformity ($\rho$) and coverage uniformity per robot; and sweep completion time ($T$), in time steps.